Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability problem

Expert replies
by dghosh2602 » Sun May 23, 2010 7:41 pm
A gambler rolls three fair 6 sided dice. What is the probability that 2 of the dice show the same number but the third shows a different number?

I don't see the flaw in my reasoning:

Lets say the gambler played the dice the first time: 1/6 is the probability for any number to come up, then for the second to match there is 1/6 again and the third not to match, its 5/6.

So the probability for the first chain of events becomes 1/6 * 1/6 * 5/6.

Now to figure out how many such chains exist:
Either the 1st and 3rd are same or 1st and the 2nd are the same or 2nd and the 3rd are the same, which gives us 3 possibilities.
This is also confirmed by the distinguishability logic: So if we are looking for all possibilities of arranging the alphabets: S, S and D => its 3! / 2! = 3.

So the total probability becomes: 1/6 * 1/6 * 5/6 * 3. But the answer is 15/36.

I don't understand the flaw in this reasoning. Please help! Thanks...
Join the discussion
Source: — Problem Solving |

by liferocks » Sun May 23, 2010 7:47 pm
your logic is almost correct,just one point.two of the dices has same number ,so probability for the second and third dice depends on the first dice face

first dice can have any value,second one has to have same value as the first so 1/6 and third has to have different value than the first so 5/6

total 1/6*5/6=5/6

now as you have said,there will be 3C2=3 ways of piling the same number dice pair.So total probability is 3*5/36=15/36

btw,if it has been mentioned that what is the probability that two dice have value '6' and the other something else ,then your ans would have been correct.Hope this will help.
“If you don't know where you are going, any road will get you there.”
Lewis Carroll
Join the discussion

by dghosh2602 » Sun May 23, 2010 8:06 pm
liferocks wrote:your logic is almost correct,just one point.two of the dices has same number ,so probability for the second and third dice depends on the first dice face

first dice can have any value,second one has to have same value as the first so 1/6 and third has to have different value than the first so 5/6

total 1/6*5/6=5/6

now as you have said,there will be 3C2=3 ways of piling the same number dice pair.So total probability is 3*5/36=15/36

btw,if it has been mentioned that what is the probability that two dice have value '6' and the other something else ,then your ans would have been correct.Hope this will help.
Got it! Your last statement helped. Thats when I realized the difference between the 2 scenarios...Thanks!
Join the discussion

by gmatjedi » Mon May 24, 2010 3:33 am
prob of rolling any number*prob of rolling same number as first number*prob of rolling different number*number of arrangements=

6/6*1/6*5/6*3=
15/36
Join the discussion